# Risk-Aware Real-Time Task Allocation for

### v1

@@ V Case Study
In addition to global and local specifications investigated in this paper, the scenario also allows for specifications explicitly assigned to individual agents, such as routing a specific bus to a designated region. The overall dynamically allocated specifications in this scenario are listed as follows.
  • When initialized, each bus X starts from T-X and is assigned a safety specification $\varphi_{(0,X)}\!:=\!\square_{[0,N]}(\mu_{\mathrm{BOX}}\wedge\lnot\mu_{\mathrm{B}\textbf{-}1}\wedge\lnot\mu_{\mathrm{B}\textbf{-}2})$.
  • At time $k=2$, a global task is given to the buses, asking them to transport tourists from GP-I, II, IV to ULP within 15 time steps. We decompose the global specification into three local specifications given by:
@@
  • At time $k=8$, a global task is given to the buses, asking them to transport tourists from GP-II, III to ULP within 15 time steps. We decompose the global specification into two local specifications given by:
@@
  • At time $k=15$, each bus $X$ is assigned with a ‘back to terminal’ specification $\varphi_{(15,X)}\!:=\!\lozenge_{[15,23]}\square_{[0,2]}\mu_{\mathrm{T}\textbf{-}\mathrm{X}}$.
All four buses have the same dynamic models described by system (1) with $A\!=\!B\!=\!I_{2}$ and $w^{i}(k)\sim\mathcal{N}(0,\sigma I_{2})$ for all $i\in\mathcal{I}$, with $\sigma=0.001$. The states and the control inputs are their positions and velocities, respectively. To incorporate the heterogeneity of practical systems, we assume that the buses have different speed limits, i.e., 444\,4m/s, 555\,5m/s, 666\,6m/s, and 777\,7m/s for buses A to D, respectively. We consider a stabilizing feedback gain $K\approx-0.618I_{2}$ for Eq. (4b), a maximal risk $r_{*,\max}=0.5$ for any global specification or explicit specification ($k=0$, $k=15$), individual maximal risk based on Eq. (11) for any local specification ($k=2$, $k=8$) and a cost $J(\bm{z},\bm{v},\bm{r})=u^{T}Ru+\sum_{i=1}^{N}r(i)$ for Eq. (9a) with $R=0.001I_{2}$. A coordinate transformation ensures all PRS are spherical and the dynamics are normalized.
@@
After decomposing, filtering, control synthesis, and auctioning, we obtain the results as seen in Figure 5.
Figure 5: The bus trajectories.
The logging information of the experiment indicates that: 1) at time $k=2$, buses A, B, and D are assigned with $\varphi_{(2,\mathrm{II})}$, $\varphi_{(2,\mathrm{IV})}$, and $\varphi_{(2,\mathrm{I})}$, respectively; 2) at time $k=8$, buses A and D are assigned with $\varphi_{(8,\mathrm{II})}$ and $\varphi_{(8,\mathrm{III})}$, respectively; 3) at time $k=15$, buses B, C, and D have accepted the return specification, while bus A has rejected it. Bus C is not assigned any task throughout the experiment and, as a result, stays at its terminal. Since buses A and D accept additional tasks at $k=8$, our method allows agents to adjust their behaviours according to newly assigned tasks. As mentioned, bus A has rejected the return task at time $k=15$. This is due to the incapability to move long distances considering its small speed limit. Additionally, all buses are confined within the ‘BOX’ region and kept outside the buildings, implying the compatibility of our method to explicitly assigned specifications. Part of the trajectory of bus D intersects with B-2 due to naive linear interpolation. In practice, critical safety can be guaranteed using nonlinear interpolation which may generate curved trajectories navigating around the building.
Remark 2
VI Conclusion
This paper considered real-time task allocation and control synthesis for stochastic heterogeneous linear multi-agent systems with probabilistic satisfaction of signal temporal logic specifications. The proposed method first decomposed the specification onto the single-agent level, to next use robustness of STL to filter the most promising agent-specification pairs. Subsequently, tube-based model predictive control was used to synthesize a control strategy and determine the local risk value for each remaining agent-specification pair. Finally, an auctioning algorithm based on the local risk values determined the optimal assignment. The proposed method was demonstrated through a multi-bus scenario.
In future work, we will consider resolving conflict due to the rejection of specifications and consider a fully centralized control synthesis. Additionally, improvements upon the heuristic approach and the conservativeness due to union bound argument will be considered.

### v2

@@ V-B Specification Decomposition: Illustrative example
Figure 5: Polytopes $H$, $B_{\mathrm{in}}(H)$, $B_{\mathrm{out}}(H)$, for Example 2-.
V-C Case Study: Technical Details
We select a control horizon $N=25$ and use a polytopic predicate $\mu_{\mathrm{R}}:=(x_{k}\in\mathrm{R})$ to denote the position $x_{k}$ of a shuttle residing within a rectangular region $\mathrm{R}$. The overall task specifications are listed as follows.
  • Initially, each shuttle $X$ should satisfy a safety specification $\varphi_{(0,X)}\!:=\!\square_{[0,N]}(\lnot\mu_{\mathrm{B}\textbf{-}1}\wedge\lnot\mu_{\mathrm{B}\textbf{-}2})$.
  • A global task is assigned at $k=2$: the shuttles should deliver tourists from GP-I, II, IV to ULP within 15 time steps. We decompose it into three local specifications:
@@ V-C Case Study: Technical Details
  • A global task is assigned at $k=8$: the shuttles should deliver tourists from GP-II, III to ULP within 15 time steps. We decompose it into two local specifications:
@@
  • At time $k=15$, each shuttle $X$ is assigned with a ‘back to terminal’ specification $\varphi_{(15,X)}\!:=\!\lozenge_{[15,23]}\square_{[0,2]}\mu_{\mathrm{T}\textbf{-}\mathrm{X}}$.
All four buses have the same dynamic models described by system (1) with $A\!=\!B\!=\!I_{2}$ and $w^{i}(k)\sim\mathcal{N}(0,\sigma I_{2})$ for all $i\in\mathcal{I}$, with $\sigma=0.001$. The states and the control inputs are their positions and velocities, respectively. To incorporate the heterogeneity of practical systems, we assume that the buses have different speed limits, i.e., 444\,4m/s, 555\,5m/s, 666\,6m/s, and 777\,7m/s for buses ‘A’ to ‘D’, respectively. We consider a stabilizing feedback gain $K\approx-0.618I_{2}$ for Eq. (4b), a maximal risk $r_{*,\max}=0.5$ for any global specification or explicit specification ($k=0$, $k=15$), individual maximal risk based on $r_{\phi^{j},\max}=\frac{r_{\phi,\max}}{\nu}$ for any local specification ($k=2$, $k=8$) and a cost $J(\bm{z},\bm{v},\bm{r})=u^{T}Ru+\sum_{i=1}^{N}r(i)$ for Eq. (5a) with $R=0.001I_{2}$.